If a binomial distribution applies with a sample size of n=20, find the values below.

a.

The probability of 55 successes if the probability of a success is 0.50

b.

The probability of at least 77 successes if the probability of a success is 0.25

c.

The expected​ value, n=20, p=0.30

d.

The standard​ deviation, n=20, p=0.30
Answer #1

a) The probability of 5 successes if the probability of a success is 0.50

P(X=5)= 0.0147857666

b) P(X>=7)=0.94234085083

c) E(X)= np= 20*0.3= 6

d) V(X)= npq= 20*0.3*0.7= 4.2

S.D= sqrt(4.2)= 2.05

NOTE: PLEASE CHECK 55 AND 77 IS TYPO error